Calculate the (regular) payback for a project with the following projected cash flows:

Year Cash Flow  
0 -2,500
1 600
2 1300
3 766
4 600

Assume the risk-adjusted WACC is 8.6%.

Enter your answer as a number with 2 decimal places of precision (i.e. 1.23)
Answer #1

Payback period is the number of years it takes to re-earn the invested amount

Invested Amount = $2,500

Cash flow in year1 = $600.

Cash flow in year2 = $1300. Cumulative cash flow in year2 = $600 + $1300 = $1900

Cash flow in year3 = $766. Cumulative cash flow in year2 = $1900 + $766 = $2666. This becomes more than the amount required. In order to complete payback, you only needed 2500 - 1900 = $600 mil.

So fraction of year 3 = $600/$766 = 0.78 year

Payback = 2 + 0.78 years = 2.78 years